1.What is the return on assets (ROA) for a firm that has a debt ratio of 0.65, a net profit margin of 6.5%, sales of $740,000, and a total asset turnover of 4?

2. Thompson Corp has a NPM of 11.5%. They reported earnings after tax of $632,500. If they had net fixed assets of $1,500,000 and current assets of $750,000, what are their total asset turnover ratios?
Answer #1
1.
Net profit margin = Net profit / Sales
6.5% = Net profit / 740000
Net profit = 740000 * 6.5% 48100
Total asset turnover = Sales / Total asset
4 = 740000 / Total assets
Total assets = 740000 / 4 185000
Return on assets (ROA) = Net profit / Total asset = 48100 / 185000 26%
2.
NPM = Earnings after tax / Sales
11.5% = 632500 / Sales
Sales = 632500 / 11.5% 5500000
Total assets = Net fixed assets + Current assets = 1500000 + 750000 2250000
Total asset turnover = Sales / Total assets = 5500000 / 2250000 2.44